my daughter has homework but I can not help her, the question is to find the rule of the following input and output;
INPUT OUTPUT
1 36
2 33
3 30
4 27
5 24
please answer this my daughter will be very happy and i will tell her all about you!

Found 2 solutions by khwang, usyim88hk:
Answer by khwang(438)


Note that the OUTPUT is a decreasing sequence and is
an arthmetic sequence with common difference -3.
The rule is : 36 - 3(INPUT -1) or 39 - 3*INPUT
Recall: The nth term of an arthmetic sequence with leading term a and
common difference d is a +(n-1)d
Kenny

Answer by usyim88hk(157)

Output = 39 - (Input)*3
(Output is equal to 39 minus triple the input)
-----------------------------------------------
when input is one
Output = 39 - (1)*3
Output = 39 - 3
Output = 36
-----------------------------------------------
when input is two
Output = 39 - (2)*3
Output = 39 - 6
Output = 33
-----------------------------------------------
So that's how it works